### v2

# Environmental impacts of astronomical research infrastructures

Abstract
Human activities degrade the Earth environment at an unprecedented scale and pace, threatening Earth-system stability, resilience and life-support functions. We can of course deny the facts, get angry about them, or try to bargain. Or we may overcome these stages of grief and move towards accepting that human activities need to change, including our own ones. The purpose of this paper is to support astronomers in this transition, by providing insights into the origins of environmental impacts in astronomical research and proposing changes that would make the field sustainable. The paper focuses on the environmental impacts of research infrastructures, since these are the dominant sources of greenhouse gas emissions in astronomy, acknowledging that impact reductions in other areas, for example professional air travelling, need also to be achieved.
keywords:
Astronomy, Research Infrastructures, Sustainability, Environment, Life Cycle Assessment1 Introduction
Carbon footprint estimates suggest that research infrastructures are the largest contributor to the greenhouse gas emissions of astronomical research (Martin et al. 2022; Knödlseder et al. 2022). What is called here research infrastructures are the astronomical observatories built on Earth and the telescopes and probes that are sent to space and that we use to learn about the Univers. Recent research suggests that the annual greenhouse gas emissions of astronomical facilities reached a value of 1.3 MtCO2e yr-1 in 2022, of which 84% stem from space missions and 16% from ground-based observatories (Knödlseder et al. 2024). The unit of MtCO2e refers here to emissions that are equivalent to one million metric tons of CO2, which is the usual unit that is used to aggregate different greenhouse gases into a single number. The research also revealed that the number of operating facilities has been constantly growing, with an average growth rate of 3.2% per year for space facilities and 1% per year for ground-based facilities over the past 30 years (Knödlseder et al. 2024). The question hence arises of whether these growth rates are environmentally sustainable.
Much of the recent research has focused on quantifying greenhouse gas emissions of astronomical research that contribute to climate change, yet human activities have additional detrimental environmental impacts that need also to be assessed and reduced. Earth system science has identified nine processes and associated boundaries that, if crossed, could generate unacceptable environmental change that threatens Earth-system stability, resilience and life-support functions. Six of these nine planetary boundaries have already been breached, including climate change, rate of biodiversity loss, interference with the nitrogen and phosphorus cycles, global freshwater use, change in land use, and chemical pollutions (Caesar et al. 2024). Recent research suggests that with ocean acidification a seventh planetary boundary has already been crossed, leaving stratospheric ozone depletion and atmospheric aerosol loading as the only processes that are still within the safe operating space (Findlay et al. 2025).
In order to gain insights into all environmental impacts of astronomical facilities we have conducted Life Cycle Assessments (LCAs) for a space system and a ground-based facility (Barret et al. 2024; Dos Santos Ilha et al. 2024). Furthermore, we have investigated how greenhouse gas emissions from astronomical facilities may evolve in the future, depending on the choices we make as a community and of funding agencies (Knödlseder et al. 2024). In this paper we summarise the results of these studies and discuss their implications on the sustainability of our research activities.
2 Estimating and reducing environmental impacts using Life Cycle Assessment and eco-design
2.1 Development and construction of X-IFU instrument
The first LCA we made concerns the X-ray Integral Field Unit (X-IFU), a cryogenic instrument that will provide from 2035 on high-resolution X-ray spectroscopy aboard the Athena space observatory of the European Space Agency (ESA) (Barret et al. 2024). The weight of the considered instrument is 221 kg, LCA results are summarised in Fig. 1. The most significant environmental impacts are “Climate change” and “Resource use”, including fossil, mineral and metal resources. Total greenhouse gas emissions amount to tCO2e with important contributions from assembly, integration and testing (driven by energy consumption of clean rooms), office work (driven by energy consumption of office buildings and staff commuting), and travelling (driven by air travels for project meetings). As most energy is produced from fossil resources their use shows a similar pattern. The most important contribution to the use of mineral and metal resources comes from the construction of pre-flight and flight hardware, driven by the use of gold in the integrated circuits of the electronics. The contribution from transport of system elements is negligible.
2.2 Construction and operations of CTAO mid-sized telescope
The second LCA we made concerns the construction, upgrade and 30 years of operations of one Mid-Sized Telescope (MST) of the Cherenkov Telescope Array Observatory (CTAO) that will be installed in the coming years on the island of La Palma, Spain (Dos Santos Ilha et al. 2024). The telescope consists of foundations (464 t), a structure including the mirrors (82 t), and the NectarCAM camera (2.1 t; all weights are in metric tons). LCA results are summarised in the left panel of Fig. 2. The most significant environmental impacts are again “Climate change” and use of mineral and metal resources. Total greenhouse gas emissions amount to tCO2e with important contributions from the construction of the telescope structure (driven by the production and processing of steel and aluminium) and the powering of the telescope during the 30 years of operations (driven by electricity generation on La Palma using diesel generators). The most important contribution to the use of mineral and metal resources comes from camera construction (driven by the use of gold in integrated circuits), and to a lesser extent from the construction of the telescope structure. Construction of the telescope foundations, transport of the telescope on site, telescope deployment and maintenance contribute little to the environmental impacts.
2.3 Reducing environmental impacts using eco-design
Using the LCA model of the CTAO mid-sized telescope we investigated whether environmental impacts can be reduced by making changes to the design or operations of the telescope. Replacing for example the steel-made counterweights by concrete blocks of the same weight would reduce the environmental impact (as measured by the single score) by 3%, with main reductions achieved on climate change and fossil resource use. Using metal machining instead of metal casting for the production of the head and yokes of the telescope positioner would reduce the single score by 12%, with impact reductions between 5% and 25%, depending on impact category. Specifically, the impact on climate change would be reduced by 14%, corresponding to greenhouse gas emission reductions of 381 tCO2e. The most dramatic impact reductions would be achieved if the diesel generators for electricity generation on La Palma were replaced by a wind turbine powered hydro storage system alike the one that is currently operating on the neighbouring island of El Hierro (see right panel of Fig. 2). While the construction of such a system would increase the environmental impact of mineral resource use by 1%, all other impacts would be reduced, with impact reductions of up to 54% on some categories. The single score would be reduced by 31%, and greenhouse gas emissions would be lowered by 46%, corresponding to tCO2e.
3 Choosing the greenhouse gas emission pathway
We studied the past evolution of the annual greenhouse gas emissions from astronomical research infrastructures based on an inventory of 586 ground-based facilities and 625 space missions, and investigated how these emissions may evolve in the future based on possible policy choices (see Fig. 3 and Knödlseder et al. 2024). Our work shows that if we choose to continue doing research as usual, emissions may either surge, or in the best case remain roughly constant, owing to an assumed (and so far observed) industrial decarbonisation rate of 1.4% per year. We could however alternatively choose to increase the decarbonisation rate, for example by asking funding agencies to make substantial investments in the decarbonisation of existing facilities. Assuming that “deep decarbonisation” levels of 3% per year can be achieved, annual greenhouse gas emissions would drop, yet not at the pace required to meet the Paris agreement targets. Alternatively we could choose to “freeze” the number of operating facilities to the current number, which would give about the same result as the “deep decarbonisation”. Combining both policy options, which probably happens naturally if we choose to divert funding from construction of new facilities to decarbonisation of existing facilities, would even more reduce the annual greenhouse gas emissions, yet still not at a rate that will halt global warming. Only if we choose to degrow the number of operating facilities by 3% per year and, at the same time, choose to invest in “deep decarbonisation”, astronomical research would be on track towards sustainability.
4 Towards sustainable astronomy
Our research suggests that eco-design of astronomical facilities may reduce their environmental impacts, yet these reductions will probably not exceed a few at best. Given the current annual growth rate of 1–3% in the number of facilities, environmental impact reductions from eco-design will be quickly outpaced by the growth of the field. As we demonstrated, without deliberately limiting the number of facilities and at the same time investing in the decarbonisation of the field, astronomy will not become sustainable. This does not imply that astronomical research will have to stop, nor that no new facilities could be built, but it implies that a more frugal approach needs to be developed.
How could astronomical research look like in the future? There would be less competition among nations, regions, and scientists and more collaboration to make astronomy a common good instead of being a reaping contest. This would imply less funding needs for new facilities and free resources for the decarbonisation of existing facilities, which would become the priority. Research would be mainly made using observations from existing facilities and archival data, and slow science would be valued instead of citation counts. Environmental impact budgets would be defined that diminish with time, making sure that astronomy stays on track on its way towards sustainability. New facilities would only be built if existing datasets in related scientific fields have been fully exploited and if their realisation does not breach the environmental impact budgets. Through these changes we would illustrate how humanity can make peace with nature, making astronomy once again a cultural good and no longer a showpiece of technological power.